OpenStudy (anonymous):

I need help on simplifying algebraic raitios. Heres one example,Simplify . Idk if you can see it though... x^2+8x/x^2

OpenStudy (anonymous):

x^2 + 8x = x(x+8) (x(x+8))/(x(x)) = (x+8)/x (x+8)/x

OpenStudy (anonymous):

Thank you, but i still dont understand how you got it. Did you just multiply?